OpenStudy (anonymous):

perform the indicated operation and, if possible, simplify 4/9+5/12

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (dumbcow):

add fractions by finding common denominator LCM of 9 and 12 is 36 4/9 = 16/36 by multiplying by 4/4 5/12 = 15/36 by multiplying by 3/3 16/36 + 15/36 = 31/36